### v1

Abstract
In view of the recent observation of $CP$-violation in bottom baryon decays, we explore $CP$-asymmetry relations in charmless four-body decays of bottom baryons featuring fully charged hadronic final states. Starting with a general effective Hamiltonian, we derive amplitude relations among various decay modes within the framework of $U$-spin symmetry. These relations lead to symmetry-based predictions for $CP$-violating rate asymmetries in $U$-spin conjugate decay channels. Furthermore, we demonstrate that all leading-order amplitude relations are violated once $U$-spin breaking effects are taken into account.
I Introduction
The discovery of $CP$ violation (CPV) in bottom baryon decays at LHCb [1] heralds a new era in experimental observations of $CP$ violation within the Standard Model. Since the observed matter-antimatter asymmetry [2] of the universe is closely tied to baryons, which make up the visible universe, $CP$ violation in baryons provides a direct connection to this phenomenon. While $CP$ violation in meson mixing and decays has been well-established through experiments [3, 4, 5, 6, 7, 8, 9], the status of CPV in baryons remained uncertain, despite significant efforts from LHCb [10, 11, 12, 13, 14, 15, 16, 17] in recent years. In conjunction with the evidence of CPV in the $\Lambda^{0}_{b}\to\Lambda K^{+}K^{-}$ decay [18], the measured $CP$ asymmetry in the $\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{+}\pi^{-}$ decay [1],
$A_{CP}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{+}\pi^{-})=(2.45\pm 0.46\pm 0.10)\%$ (1)
represents a seminal result in our understanding of the origin of $CP$ violation within the Standard Model. Since CPV in both mesons and baryons arises from the CKM mechanism [19, 20] that governs quark transitions, a crucial test of the theory comes from flavor symmetries, particularly given the challenges in making precise theoretical predictions for $CP$ violation in multibody baryon decays. Several approaches have been proposed to study $CP$ violation in multibody bottom baryon decays [21, 22, 23], including a method to extract the $CP$-violating weak phase in charmless $b$-baryon decays [24, 25].
In this paper, we use a subclass of flavor symmetry, known as the $U$-spin symmetry [26, 27], reflecting the symmetry between down type quarks $d$ and $s$ to relate $b\to d$ and $b\to s$ transition amplitudes [28, 29]. In our approach, we first construct the $U$-spin decomposition [28, 30, 29, 31] of the most general Hamiltonian mediating a transition from the bottom baryon to a four body charmless final state. In the next step, we consider the dimension-6 effective Hamiltonian of the SM that mediates such transition and map its $U$-spin components to the corresponding components of the general Hamiltonian. At the leading-order $U$-spin limit, a set of amplitude relations among the four-body final states accessible to the bottom baryons naturally emerges. Moreover, this method allows us to extend the analysis beyond the exact $U$-spin symmetry limit and determine whether specific leading-order amplitude relations continue to hold. This feature is particularly significant given that $U$-spin symmetry is broken by the mass difference between the $d$ and $s$ quarks, with the breaking estimated to be at the level of $\sim$ 30% [29]. We wish to emphasize, however, that our $U$-spin symmetry–based argument is intended primarily as a guiding framework within a data-driven approach, useful for estimating decay amplitudes and identifying promising four-body bottom baryon decays where sizable $CP$ violation may be observed.
There are several $U$-spin based analysis [32, 33, 34, 35] available that connects the recently measured CPV observed in $\Lambda^{0}_{b}$ decay with their corresponding $U$-spin partners under specific assumptions. Our approach differs from previous analyses by first enumerating the maximum number of invariant $U$-spin amplitudes permitted in the $B_{b}\to B_{1}PP^{\prime}P_{1}$ transition, where $B_{b}$ is the initial bottom baryon, $P$ and $P^{\prime}$ are pseudoscalar mesons with the same electric charge, and $P_{1}$ is a pseudoscalar meson with opposite charge, and then systematically relating or neglecting these amplitudes based on the $U$-spin decomposition of the Standard Model dimension-6 effective Hamiltonian. It is worth highlighting that our approach closely mirrors the method first introduced in [36] to study $B$-meson decays using $SU(3)$ flavor symmetry, and subsequently applied to two-body hadronic decays of bottom baryons in [37, 38].
In Section (II), we discuss the theoretical framework for performing a $U$-spin decomposition of bottom baryon decay amplitudes involving fully charged four-body final states. Section (III) focuses on derivation of symmetry-based relations between physical observables, such as $CP$-violating rate asymmetries. We conclude by commenting on the effect of $U$-spin breaking on these observables.
II Application of $U$-spin to decay amplitudes
Under $U$-spin, the final state charged pseudoscalar mesons form doublets of $U$-spin [39, 40] expressed as,
$P,\,P^{\prime}=\begin{pmatrix}\pi^{-}\\ K^{-}\end{pmatrix}=\begin{pmatrix}-|\frac{1}{2},\frac{1}{2}\rangle\\ -|\frac{1}{2},-\frac{1}{2}\rangle\end{pmatrix},\quad\begin{pmatrix}K^{+}\\ \pi^{+}\end{pmatrix}=\begin{pmatrix}|\frac{1}{2},\frac{1}{2}\rangle\\ -|\frac{1}{2},-\frac{1}{2}\rangle\end{pmatrix}$ (2)
Similarly, a subset of the light baryons belonging to the octet under $SU(3)$-flavor form distinct doublets under $U$-spin given below,
$B_{1}^{+}=\begin{pmatrix}p^{+}\\ \Sigma^{+}\end{pmatrix}=\begin{pmatrix}|\frac{1}{2},\frac{1}{2}\rangle\\ |\frac{1}{2},-\frac{1}{2}\rangle\end{pmatrix},\quad B_{1}^{-}=\begin{pmatrix}\Sigma^{-}\\ \Xi^{-}\end{pmatrix}=\begin{pmatrix}|\frac{1}{2},\frac{1}{2}\rangle\\ |\frac{1}{2},-\frac{1}{2}\rangle\end{pmatrix}.$ (3)
The bottom baryon under $SU(3)$-flavor symmetry transforms as an anti-triplet while its neutral components $\Lambda^{0}_{b}$ and $\Xi_{b}^{0}$ form an $U$-spin doublet $B_{b}=\begin{pmatrix}\Lambda^{0}_{b}&\Xi_{b}^{0}.\end{pmatrix}$ Given the initial state ($\mathbf{i}$) and final states ($\mathbf{f}$), we can write down the most general transition connecting the initial and final state symbolically as,
$\mathcal{H}_{\text{eff}}\equiv\mathbf{f}\otimes\overline{\mathbf{i}}.$ (4)
For the final state we use the order $(P\otimes P^{{}^{\prime}})\otimes P_{1}\otimes B_{1}$ to calculate the tensor product and subsequently apply the Wigner-Eckart theorem for $U$-spin decomposition. The final state $\mathbf{f}$ made off two identically charged mesons, a third oppositely charged meson and a light baryon must be symmetrized over the identically charged mesons tensor product state as they are indistinguishable under exact $U$-spin [24]. Operationally we have,
$\{P^{\pm}P^{{}^{\prime}\pm}\}=\frac{1}{\sqrt{2}}(P^{\pm}P^{\prime\pm}+P^{{}^{\prime}\pm}P^{\pm})$ (5)
We highlight that this symmetrization process misses an additional $\sqrt{2}$ factor whenever two identical mesons $P^{\pm}(U,U_{3})=P^{{}^{\prime}\pm}(U^{{}^{\prime}}=U,U_{3}^{{}^{\prime}}=U_{3})$ are encountered which we account for in our calculation. Now, we can perform the $U$-spin decomposition of the decay amplitude as,
$\mathcal{A}(B_{b}\to B_{1}^{\pm}PP^{\prime}P_{1})=\quad\sum^{U^{P}+U^{P^{\prime}},U^{M^{0}}+U^{P_{1}}\dots}_{\mathclap{\begin{subarray}{c}\{U^{M^{0}}=|U^{P}-U^{P^{\prime}}|,\,U^{M}=|U^{M^{0}}-U^{P_{1}}|,\,U^{F}\dots\}\\ U^{P}_{3}+U^{P^{\prime}}_{3}=U_{3}^{M_{0}},\,U^{M_{0}}_{3}+U_{3}^{P_{1}}=U_{3}^{M}\\ U_{3}^{M}+U_{3}^{B_{1}}=U_{3}^{F},\,U_{3}^{F}-U_{3}^{B_{b}}=U_{3}^{\mathcal{H}}\end{subarray}}}C^{\{U^{P}_{3}U^{P^{\prime}}_{3}U_{3}^{M_{0}}\}}_{U^{P}U^{P^{\prime}}U^{M_{0}}}C^{U_{3}^{M_{0}}\,U_{3}^{P_{1}}\,U_{3}^{M}}_{U^{M^{0}}\,U^{P_{1}}\,U^{M}}$ (6)
×$\times C^{U_{3}^{M}\,U_{3}^{B_{1}}\,U_{3}^{F}}_{U^{M}\,U^{B_{1}}\,U^{F}}C^{U_{3}^{F}\,(-U_{3}^{B_{b}})\,U_{3}^{\mathcal{H}}}_{U^{F}\,U^{B_{b}}\,U^{\mathcal{H}}}A^{U^{M},U^{F},\pm}(U^{\mathcal{H}})$
where $A^{i,j,\pm}(U^{\mathcal{H}})$ denotes the $U$-spin subamplitudes and $C^{U_{3}^{a}U_{3}^{b}U_{3}^{c}}_{U^{a}U^{b}U^{c}}$ represents the $SU(2)$ Clebsch–Gordan coefficients. The $\pm$ label in $A^{i,j,\pm}(U^{\mathcal{H}})$ distinguishes the $U$-spin doublet to which the final state light baryon belongs, and in general $A^{i,j,+}(U^{\mathcal{H}})$ and $A^{i,j,-}(U^{\mathcal{H}})$ are unrelated. The subamplitudes $A^{i,j,\pm}(U^{\mathcal{H}})$ in the most general case also involve unknown dynamical coefficients, for example, in the form of Wilson coefficients and CKM factors, that only get fixed when a particular form of the effective Hamiltonian is assumed. Therefore, apriori $A^{i,j,\pm}(U^{\mathcal{H}})$ are all independent of each other. We now proceed to a discussion of the lowest order effective Hamiltonian [41, 42, 43] mediating charmless bottom baryon $\Delta|S|=1$ and $\Delta S=0$ transitions. Both parts of the effective Hamiltonian is composed of the operators $O_{1}$,…, $O_{10}$ and can be written as:
$\mathcal{H}_{\text{eff}}=\frac{4G_{F}}{\sqrt{2}}\Big{[}\lambda^{(s)}_{u}\Big{(}C_{1}(O^{(u)}_{1}-O^{(c)}_{1})+C_{2}(O^{(u)}_{2}-O^{(c)}_{2})\Big{)}-\lambda^{(s)}_{t}\sum_{i=1,2}C_{i}O^{(c)}_{i}-\lambda^{(s)}_{t}\sum_{i=3}^{10}C_{i}O_{i}^{(s)}$
$+\lambda^{(d)}_{u}\Big{(}C_{1}(O^{(u)}_{1}-O^{(c)}_{1})+C_{2}(O^{(u)}_{2}-O^{(c)}_{2})\Big{)}-\lambda^{(d)}_{t}\sum_{i=1,2}C_{i}O^{(c)}_{i}-\lambda^{(d)}_{t}\sum_{i=3}^{10}C_{i}O_{i}^{(d)}\Big{]},$ (7)
where $V_{ub}V_{us}^{*}=\lambda_{u}^{s}$, $V_{ub}V_{ud}^{*}=\lambda_{u}^{d}$, $V_{tb}V_{ts}^{*}=\lambda_{t}^{s}$, $V_{tb}V_{td}^{*}=\lambda_{t}^{d}$ are the CKM elements and $C_{i}$ s are the Wilson coefficients. $O_{1}$ and $O_{2}$ are the “tree” operators:
$O_{1}^{(u)}$ $=(\overline{u}^{i}_{L}\gamma^{\mu}b^{j}_{L})(\overline{s}^{j}_{L}\gamma_{\mu}u^{i}_{L})$
$O_{1}^{(c)}$ $=(\overline{c}^{i}_{L}\gamma^{\mu}b^{j}_{L})(\overline{s}^{j}_{L}\gamma_{\mu}c^{i}_{L})$
$O_{2}^{(u)}$ $=(\overline{u}^{i}_{L}\gamma^{\mu}b^{i}_{L})(\overline{s}^{j}_{L}\gamma_{\mu}u^{j}_{L})$
$O_{2}^{(c)}$ $=(\overline{c}^{i}_{L}\gamma^{\mu}b^{i}_{L})(\overline{s}^{j}_{L}\gamma_{\mu}c^{j}_{L}).$ (8)
$O_{3}$, …, $O_{6}$ are the “gluonic penguin” operators:
$O_{3}^{(s)}$ $=(\overline{s}^{i}_{L}\gamma^{\mu}b^{i}_{L})\sum_{q=u,d,s}(\overline{q}^{j}_{L}\gamma_{\mu}q^{j}_{L})$
$O_{4}^{(s)}$ $=(\overline{s}^{i}_{L}\gamma^{\mu}b^{j}_{L})\sum_{q=u,d,s}(\overline{q}^{j}_{L}\gamma_{\mu}q^{i}_{L})$
$O_{5}^{(s)}$ $=(\overline{s}^{i}_{L}\gamma^{\mu}b^{i}_{L})\sum_{q=u,d,s}(\overline{q}^{j}_{R}\gamma_{\mu}q^{j}_{R})$
$O_{6}^{(s)}$ $=(\overline{s}^{i}_{L}\gamma^{\mu}b^{j}_{L})\sum_{q=u,d,s}(\overline{q}^{j}_{R}\gamma_{\mu}q^{i}_{R}).$ (9)
and finally $O_{7}$, …, $O_{10}$ are the four “Electroweak penguin” operators:
$O_{7}^{(s)}$ $=\frac{3}{2}(\overline{s}^{i}_{L}\gamma^{\mu}b^{i}_{L})\sum_{q=u,d,s}e_{q}(\overline{q}^{j}_{R}\gamma_{\mu}q^{j}_{R}),$
$O_{8}^{(s)}$ $=\frac{3}{2}(\overline{s}^{i}_{L}\gamma^{\mu}b^{j}_{L})\sum_{q=u,d,s}e_{q}(\overline{q}^{j}_{R}\gamma_{\mu}q^{i}_{R}),$ (10)
$O_{9}^{(s)}$ $=\frac{3}{2}(\overline{s}^{i}_{L}\gamma^{\mu}b^{i}_{L})\sum_{q=u,d,s}e_{q}(\overline{q}^{j}_{L}\gamma_{\mu}q^{j}_{L}),$
$O_{10}^{(s)}$ $=\frac{3}{2}(\overline{s}^{i}_{L}\gamma^{\mu}b^{j}_{L})\sum_{q=u,d,s}e_{q}(\overline{q}^{j}_{L}\gamma_{\mu}q^{i}_{L}).$ (11)
Restricting ourselves to a discussion of $U$-spin of these operators, we find that the tree operators $O_{1}$ and $O_{2}$ transform as a $U$-spin doublet with the up component corresponding to $b\to d$ transition ($\Delta S=0$) while the down component corresponds to $b\to s$ transition ($\Delta|S|=1$) transition. The relevant CKM elements which form part of the dynamical coefficients are given, respectively, as $V_{ub}V_{ud}^{*}$ and $V_{ub}V_{us}^{*}$. Since the combination $d\overline{d}+s\overline{s}$ transforms as a $U$-spin singlet, the complete set of QCD and Electroweak penguin four-Fermi operators ($O_{3}\dots O_{10}$) transform as distinct $U$-spin doublets. Interestingly, unlike isospin, which is explicitly broken by electromagnetic interactions, leading to isospin-violating effects in the electroweak penguin operators, no such complication arises in the case of $U$-spin, since the $d$ and $s$ quarks carry the same electric charge. Therefore, to the leading order in $U$-spin, the dim-6 SM effective Hamiltonian only induces a doublet in Eq (6). We observe that certain higher dimensional $U$-spin representations present in the general Hamiltonian have vanishing dynamical coefficients (See Appendix (A)) corresponding to non-existent $U$-spin subamplitudes.
III Amplitude Relations
Possible four body all-electrically charged final states originating from charmless bottom baryon decays are listed in Table (1).
$\Delta|S|=1$ modes $\Delta S=0$ modes $\Delta|S|=1$ modes $\Delta S=0$ modes
$\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+}$ $\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+}$ $\Xi^{0}_{b}\to p^{+}K^{-}K^{-}\pi^{+}$ $\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+}$
$\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-}$ $\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-}$ $\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}K^{-}$ $\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}\pi^{-}$
$\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+}$ $\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+}$ $\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}\pi^{+}$ $\Lambda^{0}_{b}\to p^{+}\pi^{-}K^{-}K^{+}$
$\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-}$ $\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-}$ $\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-}$ $\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}\pi^{-}$
$\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+}$ $\Xi^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+}$ $\Xi^{0}_{b}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-}$ $\Lambda^{0}_{b}\to\Sigma^{-}K^{+}K^{+}K^{-}$
$\Lambda^{0}_{b}\to\Xi^{-}K^{+}\pi^{+}\pi^{-}$ $\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}K^{+}K^{-}$ $\Xi^{0}_{b}\to\Sigma^{+}K^{-}K^{-}K^{+}$ $\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+}$
$\Lambda^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+}$ $\Xi^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+}$
$\Lambda^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-}$ $\Xi^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-}$
Table 1: Charmless four body decays of bottom baryons involving all charged final states
We choose to represent the $\Delta S=0$ and $\Delta|S|=1$ decay amplitudes and the complete set of $U$-spin subamplitudes in terms of column matrices $\mathcal{A}$ and $A$ respectively such that there is coefficient matrix ($\mathbb{C}$) that links both basis through the equation,
$\mathcal{A}=\mathbb{C}A$ (12)
The explicit form the matrix $\mathbb{C}$ is given in Appendix (A). We notice that there are in general 14 independent $U$-spin subamplitudes describing all 28 decays. Alternatively, these 28 decay modes can be grouped based on the charge of the final state baryon. In that case, 14 decay modes with a positively charged baryon are described in terms of 7 $A^{i,j,+}(U^{\mathcal{H}})$ subamplitudes while the remaining 14 decay modes with a negatively charged baryon are described in terms of 7 $A^{i,j,-}(U^{\mathcal{H}})$ subamplitudes. At this stage, no simple amplitude relation between only two decay modes can be established. Now assuming the SM Hamiltonian expressed in Eq (II), which contributes to the $U$-spin subamplitudes $A^{\frac{1}{2},0,\pm}(U^{\mathcal{H}}=\frac{1}{2}),\,A^{\frac{1}{2},1,\pm}(U^{\mathcal{H}}=\frac{1}{2})$ and $A^{\frac{3}{2},1,\pm}(U^{\mathcal{H}}=\frac{1}{2})$ but not to $A^{\frac{1}{2},1,\pm}(U^{\mathcal{H}}=\frac{3}{2}),\,A^{\frac{3}{2},1,\pm}(U^{\mathcal{H}}=\frac{3}{2}),\,A^{\frac{3}{2},2,\pm}(U^{\mathcal{H}}=\frac{3}{2}),\,A^{\frac{3}{2},2,\pm}(U^{\mathcal{H}}=\frac{5}{2})$ for both the tree and penguin operators, the number of independent $U$-spin subamplitudes can be reduced to 6, three for each class of decays with a positively or negatively charged baryon in the final state. We can further decompose the decay amplitudes in terms of tree ($\mathcal{A}_{T}$) and penguin ($\mathcal{A}_{P}$) parts by factoring out the CKM elements $\lambda^{s,d}_{u,t}$ from the decay amplitude and write,
$\mathcal{A}^{(s,d)}=\lambda_{u}^{s,d}\mathcal{A}_{T}+\lambda^{s,d}_{t}\mathcal{A}_{P}$ (13)
where $(s,d)$ denote the $\Delta|S|=1$, $\Delta S=0$ process respectively.
$B_{b}\to B_{1}^{+}PP^{{}^{\prime}}P_{1}$ modes $B_{b}\to B_{1}^{-}PP^{{}^{\prime}}P_{1}$ modes $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}K^{-}K^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}K^{+}K^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}\pi^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}\pi^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to p^{+}\pi^{-}K^{-}K^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}K^{-})$
Table 2: Amplitude $U$-spin relations between $\Delta|S|=1$ and $\Delta S=0$ four body bottom baryon decays, assuming the unbroken SM Hamiltonian. The subscript $T$ and $P$ denotes the tree and penguin part of the full decay amplitude respectively.
We obtain the following amplitude relations connecting the $\Delta|S|=1$ and $\Delta S=0$ sectors given in Table (2). $A_{CP}$ is defined subsequently as,
$A_{CP}=$ $\frac{\Gamma(B_{b}\to B_{1}PP^{\prime}P_{1})-\Gamma(\overline{B_{b}}\to\overline{B_{1}}\overline{P}\overline{P^{\prime}}\overline{P_{1}})}{\Gamma(B_{b}\to B_{1}PP^{\prime}P_{1})+\Gamma(\overline{B_{b}}\to\overline{B_{1}}\overline{P}\overline{P^{\prime}}\overline{P_{1}})}$
$=$  $\frac{-\,4\textbf{J}\times\text{Im}\Big{[}\mathcal{A}^{*}_{T}(B_{b}\to B_{1}PP^{\prime}P_{1})\mathcal{A}_{P}(B_{b}\to B_{1}PP^{\prime}P_{1})\Big{]}}{|\mathcal{A}(B_{b}\to B_{1}PP^{\prime}P_{1})|^{2}+|\mathcal{A}(\overline{B_{b}}\to\overline{B_{1}}\overline{P}\overline{P^{\prime}}\overline{P_{1}})|^{2}},$ (14)
where J is the Jarlskog invariant given by $\text{Im}(V_{ub}V_{ud}^{*}V_{tb}^{*}V_{td})=-\text{Im}(V_{ub}V_{us}^{*}V_{tb}^{*}V_{ts})=\textbf{J}$. Neglecting any $U$-spin breaking effects from phase space factors, we can relate the $CP$-asymmetry in $\Delta|S|=1$ and $\Delta S=0$ decay modes using Eq (15),
$\frac{A_{CP}(B_{bi}\to B_{1}^{a}P^{b}P^{{}^{\prime}c}P_{1}^{d})|_{\Delta S}}{A_{CP}(B_{bj}\to B_{1}^{e}P^{f}P^{{}^{\prime}g}P_{1}^{h})|_{\Delta S^{\prime}}}\!\simeq-\frac{\tau_{\mathcal{B}_{bi}}}{\tau_{\mathcal{B}_{bj}}}\frac{Br(B_{bj}\to B_{1}^{e}P^{f}P^{{}^{\prime}g}P_{1}^{h})}{Br(B_{bi}\to B_{1}^{a}P^{b}P^{{}^{\prime}c}P_{1}^{d}))},\qquad\Delta S\neq\Delta S^{\prime}$ (15)
where $i$, $j$, and $a,\,e$ $b,\,f$, $c,\,g$, $d,\,h$ indices correspond to the various baryon and mesons in the initial and final states and the negative sign accounts for the relative sign between the Jarlskog factor in $\Delta|S|=1$ and $\Delta S=0$ decay modes. These $A_{CP}$-relations can be read off from Table (2). In particular, measurements for the following branching ratios of $\Lambda^{0}_{b}$-baryon [44],
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})=(5.0\pm 0.5)\times 10^{-5}$
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}K^{+})=(4.0\pm 0.6)\times 10^{-6}$
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})=(1.25\pm 0.13)\times 10^{-5}$ (16)
$Br(\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+})=(2.08\pm 0.21)\times 10^{-5}$
can be used to express the $CP$-asymmetries in the decay modes $\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+}$, $\Xi^{0}_{b}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+}$, $\Xi^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+}$ and $\Xi^{0}_{b}\to\Sigma^{+}K^{-}K^{-}K^{+}$ respectively.
We now discuss the effect of first order breaking of $U$-spin originating form unequal masses of the $d$ an $s$ quark. $U$-spin breaking is introduced by a triplet under $U$-spin ($U_{\text{brk}}=1,U^{3}_{\text{brk}}=0$) and by taking the tensor product with the unbroken Hamiltonian that transforms as a doublet. The effect of this $U$-spin breaking is a correction to the unbroken Hamiltonian given by,
$\mathcal{H}_{\text{brk}}^{\Delta|S|=0,1}\propto\mathcal{\epsilon}(1,0)\otimes(\frac{1}{2},\pm\frac{1}{2})=\epsilon\Big{(}\sqrt{\frac{2}{3}}(\frac{3}{2},\pm\frac{1}{2})\mp\sqrt{\frac{1}{3}}(\frac{1}{2},\pm\frac{1}{2})\Big{)}$ (17)
This induces non-zero dynamical coefficients for the $U$-spin subamplitudes $A^{\frac{1}{2},1,\pm}(\frac{3}{2}),\,A^{\frac{3}{2},1,\pm}(\frac{3}{2}),$ $A^{\frac{3}{2},2,\pm}(\frac{3}{2})$ while $A^{\frac{1}{2},0}(\frac{1}{2}),\,A^{\frac{1}{2},1}(\frac{1}{2}),\,A^{\frac{3}{2},1}(\frac{1}{2})$ pick up $U$-spin breaking corrections. Collectively, we can parametrize these $U$-spin breaking amplitudes as,
$A_{\text{brk}}^{\frac{1}{2},0,\pm}(\frac{1}{2})=\frac{\epsilon_{1}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},0,\pm}(\frac{1}{2})$
$A_{\text{brk}}^{\frac{1}{2},1,\pm}(\frac{1}{2})=\frac{\epsilon_{2}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},1,\pm}(\frac{1}{2})$
$A_{\text{brk}}^{\frac{1}{2},1,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{3}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},1,\pm}(\frac{3}{2})$ (18)
$A_{\text{brk}}^{\frac{3}{2},1,\pm}(\frac{1}{2})=\frac{\epsilon_{4}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},1,\pm}(\frac{1}{2})$
$A_{\text{brk}}^{\frac{3}{2},1,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{5}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},1,\pm}(\frac{3}{2})$
$A_{\text{brk}}^{\frac{3}{2},2,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{6}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},2,\pm}(\frac{3}{2})$
The combined result of leading order unbroken Hamiltonian and first order $U$-spin breaking contribution to the decay amplitudes is given in Appendix (B). The first order $U$-spin breaking corrections have the same structures but opposite signs for the decay modes related by $U$-spin as found in [29]. This increases the rank of the coefficient matrices to 12 and the leading order relations between $U$-spin conjugate amplitude pairs no longer hold. Naturally, the leading order $CP$-asymmetry relations evident from Table (2) also breaks down once $U$-spin breaking effects are considered. Any reduction among the set of $U$-spin subamplitudes would require knowledge of their relative magnitudes, which lies beyond the scope of the present work. We note that if a single $U$-spin subamplitude dominates, the framework becomes significantly more predictive.
IV Conclusion
We have investigated charmless bottom baryon decays into fully charged four-body final states within the realm of $U$-spin symmetry. Bearing in mind that significant $U$-spin breaking is observed in $B$-meson decays, we set up a framework to analyze bottom baryon decays where $U$-spin symmetry can be arbitrarily broken. We obtained amplitude relations between decay modes and based on those relations predicted $CP$ asymmetry relations between a number of decay modes. While $U$-spin breaking effect spoils the relations between observables, the predictions offer a roadmap for probing $CP$ violation in bottom baryon decays and provide a systematic approach to test the validity of $U$-spin symmetry in future experimental measurements.
Appendix A
$U$-spin amplitude decomposition of $B_{b}\to B_{1}^{+}PP^{\prime}P_{1}$ decay modes for general Hamiltonian:
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}K^{+})\\ \end{pmatrix}=\left(\begin{array}[]{ccccccc}-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&\frac{1}{3}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&\frac{2}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&-\frac{1}{2}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&\frac{1}{3}&-\frac{1}{3}&\frac{1}{3\sqrt{2}}&-\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&\frac{2}{3}&\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{6}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{3}{2})\\ A^{\frac{3}{2},1,+}(\frac{1}{2})\\ A^{\frac{3}{2},1,+}(\frac{3}{2})\\ A^{\frac{3}{2},2,+}(\frac{3}{2})\\ A^{\frac{3}{2},2,+}(\frac{5}{2})\end{pmatrix}$ (33)
$U$-spin amplitude decomposition of $B_{b}\to B_{1}^{-}PP^{\prime}P_{1}$ decay modes for general Hamiltonian:
$\vspace{-5cm}\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}K^{+}\pi^{-})\\ \end{pmatrix}=\left(\begin{array}[]{ccccccc}\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&\frac{1}{3}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&\frac{2}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&\frac{1}{3}&-\frac{1}{3}&\frac{1}{3\sqrt{2}}&-\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&-\frac{1}{2}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&\frac{2}{3}&\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{6}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{3}{2})\\ A^{\frac{3}{2},1,-}(\frac{1}{2})\\ A^{\frac{3}{2},1,-}(\frac{3}{2})\\ A^{\frac{3}{2},2,-}(\frac{3}{2})\\ A^{\frac{3}{2},2,-}(\frac{5}{2})\end{pmatrix}$ (48)
$U$-spin amplitude decomposition of $B_{b}\to B_{1}PP^{\prime}P_{1}$ decay modes using the SM Hamiltonian,
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}K^{+}\pi^{-})\\ \end{pmatrix}=\mathcal{C}_{T(P)}\left(\begin{array}[]{cccccc}-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}\\ 0&0&0&0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}\\ 0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}\\ 0&0&\frac{1}{\sqrt{2}}&0&0&0\\ 0&0&0&0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&0&0&0\\ 0&0&0&0&0&\frac{1}{\sqrt{2}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}\\ 0&0&0&0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}\\ 0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}\\ 0&0&\frac{1}{\sqrt{2}}&0&0&0\\ 0&0&0&0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&0&0&0\\ 0&0&0&0&0&\frac{1}{\sqrt{2}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{1}{2})\\ A^{\frac{3}{2},1,+}(\frac{1}{2})\\ A^{\frac{1}{2},0,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{1}{2})\\ A^{\frac{3}{2},1,-}(\frac{1}{2})\\ \end{pmatrix}$ (77)
where $\mathcal{C}_{T}$ and $\mathcal{C}_{P}$ are the Wilson Coefficients of tree operators ($O_{1},\,O_{2}$) and penguin operators ($O_{3}\dots O_{10}$) respectively.
Appendix B
$U$-spin amplitude decomposition of $B_{b}\to B_{1}PP^{\prime}P_{1}$ decay modes incorporating first order $U$-spin breaking to the SM Hamiltonian,