### v2

# CPCP-violation and $U$-spin symmetry in four-body bottom baryon decays

Abstract
In view of the recent observation of $CP$-violation in bottom baryon decays, we explore $CP$-asymmetry relations in charmless four-body decays of bottom baryons featuring fully charged hadronic final states. Starting with a general effective Hamiltonian, we derive amplitude relations among various decay modes within the framework of $U$-spin symmetry. These relations lead to symmetry-based predictions for $CP$-violating rate asymmetries in $U$-spin conjugate decay channels. Furthermore, we demonstrate that all leading-order amplitude relations are violated once $U$-spin breaking effects are taken into account.
I Introduction
The discovery of $CP$ violation (CPV) in bottom baryon decays at LHCb [1] heralds a new era in experimental observations of $CP$ violation within the Standard Model. Since the observed matter-antimatter asymmetry [2] of the universe is closely tied to baryons, which make up the visible universe, $CP$ violation in baryons provides a direct connection to this phenomenon. While $CP$ violation in meson mixing and decays has been well-established through experiments [3, 4, 5, 6, 7, 8, 9], the status of CPV in baryons remained uncertain, despite significant efforts from LHCb [10, 11, 12, 13, 14, 15, 16, 17] in recent years. In conjunction with the evidence of CPV in the $\Lambda^{0}_{b}\to\Lambda K^{+}K^{-}$ decay [18], the measured $CP$ asymmetry in the $\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{+}\pi^{-}$ decay [1],
$A_{CP}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{+}\pi^{-})=(2.45\pm 0.46\pm 0.10)\%$ (1)
represents a seminal result in our understanding of the origin of $CP$ violation within the Standard Model. Since CPV in both mesons and baryons arises from the CKM mechanism [19, 20] that governs quark transitions, a crucial test of the theory comes from flavor symmetries, particularly given the challenges in making precise theoretical predictions for $CP$ violation in multibody baryon decays. Several approaches have been proposed to study $CP$ violation in multibody bottom baryon decays [21, 22, 23], including a method to extract the $CP$-violating weak phase in charmless bottom baryon decays [24, 25].
In this paper, we use a subclass of flavor symmetry, known as the $U$-spin symmetry [26, 27], reflecting the symmetry between down type quarks $d$ and $s$ to relate $b\to d$ and $b\to s$ transition amplitudes [28, 29]. In our approach, we first construct the $U$-spin decomposition [28, 30, 29, 31] of the most general Hamiltonian mediating a transition from the bottom baryon to a four body charmless final state. In the next step, we consider the dimension-6 effective Hamiltonian of the SM that mediates such transition and map its $U$-spin components to the corresponding components of the general Hamiltonian. At the leading-order $U$-spin limit, a set of amplitude relations among the four-body final states accessible to the bottom baryons naturally emerges. Moreover, this method allows us to extend the analysis beyond the exact $U$-spin symmetry limit and determine whether specific leading-order amplitude relations continue to hold. This feature is particularly significant given that $U$-spin symmetry is broken by the mass difference between the $d$ and $s$ quarks, with the breaking estimated to be at the level of $\sim$ 30% [29]. We wish to emphasize, however, that our $U$-spin symmetry–based argument is intended primarily as a guiding framework within a data-driven approach, useful for estimating decay amplitudes and identifying promising four-body bottom baryon decays where sizable $CP$ violation may be observed.
There are several $U$-spin based analysis [32, 33, 34, 35] available that connects the recently measured CPV observed in $\Lambda^{0}_{b}$ decay with their corresponding $U$-spin partners under specific assumptions. Our approach differs from previous analyses by first enumerating the maximum number of invariant $U$-spin amplitudes permitted in the $B_{b}\to B_{1}PP^{\prime}P_{1}$ transition, where $B_{b}$ is the initial bottom baryon, $P$ and $P^{\prime}$ are pseudoscalar mesons with the same electric charge, and $P_{1}$ is a pseudoscalar meson with opposite charge, and then systematically relating or neglecting these amplitudes based on the $U$-spin decomposition of the Standard Model dimension-6 effective Hamiltonian. It is worth highlighting that our approach closely mirrors the method first introduced in [36] to study $B$-meson decays using $SU(3)$ flavor symmetry, and subsequently applied to two-body hadronic decays of bottom baryons in [37, 38].
In Section (II), we discuss the theoretical framework for performing a $U$-spin decomposition of bottom baryon decay amplitudes involving fully charged four-body final states. Section (III) focuses on derivation of symmetry-based relations between physical observables, such as $CP$-violating rate asymmetries. We conclude by commenting on the effect of $U$-spin breaking on these observables.
II Application of $U$-spin to decay amplitudes
Under $U$-spin, the final state charged pseudoscalar mesons form doublets of $U$-spin [39, 40] expressed as,
$P,\,P^{\prime}=\begin{pmatrix}\pi^{-}\\ K^{-}\end{pmatrix}=\begin{pmatrix}-|\frac{1}{2},\frac{1}{2}\rangle\\ -|\frac{1}{2},-\frac{1}{2}\rangle\end{pmatrix},\quad\begin{pmatrix}K^{+}\\ \pi^{+}\end{pmatrix}=\begin{pmatrix}|\frac{1}{2},\frac{1}{2}\rangle\\ -|\frac{1}{2},-\frac{1}{2}\rangle\end{pmatrix}$ (2)
Similarly, a subset of the light baryons belonging to the octet under $SU(3)$-flavor form distinct doublets under $U$-spin given below,
$B_{1}^{+}=\begin{pmatrix}p^{+}\\ \Sigma^{+}\end{pmatrix}=\begin{pmatrix}|\frac{1}{2},\frac{1}{2}\rangle\\ |\frac{1}{2},-\frac{1}{2}\rangle\end{pmatrix},\quad B_{1}^{-}=\begin{pmatrix}\Sigma^{-}\\ \Xi^{-}\end{pmatrix}=\begin{pmatrix}|\frac{1}{2},\frac{1}{2}\rangle\\ |\frac{1}{2},-\frac{1}{2}\rangle\end{pmatrix}$. (3)
The bottom baryon under $SU(3)$-flavor symmetry transforms as an anti-triplet while its neutral components $\Lambda^{0}_{b}$ and $\Xi_{b}^{0}$ form an $U$-spin doublet $B_{b}=\begin{pmatrix}\Lambda^{0}_{b}&\Xi_{b}^{0}.\end{pmatrix}$ Given the initial state ($\mathbf{i}$) and final states ($\mathbf{f}$), we can write down the most general transition connecting the initial and final state symbolically as,
$\mathcal{H}_{\text{eff}}\equiv\mathbf{f}\otimes\overline{\mathbf{i}}.$ (4)
For the final state we use the order $(P\otimes P^{{}^{\prime}})\otimes P_{1}\otimes B_{1}$ to calculate the tensor product and subsequently apply the Wigner-Eckart theorem for $U$-spin decomposition. The final state $\mathbf{f}$ made off two identically charged mesons, a third oppositely charged meson and a light baryon must be symmetrized over the identically charged mesons tensor product state as they are indistinguishable under exact $U$-spin [24]. Operationally we have,
$\{P^{\pm}P^{{}^{\prime}\pm}\}=\frac{1}{\sqrt{2}}(P^{\pm}P^{\prime\pm}+P^{{}^{\prime}\pm}P^{\pm})$ (5)
We highlight that this symmetrization process misses an additional $\sqrt{2}$ factor whenever two identical mesons $P^{\pm}(U,U_{3})=P^{{}^{\prime}\pm}(U^{{}^{\prime}}=U,U_{3}^{{}^{\prime}}=U_{3})$ are encountered which we account for in our calculation. Now, we can perform the $U$-spin decomposition of the decay amplitude as,
$\mathcal{A}(B_{b}\to B_{1}^{\pm}PP^{\prime}P_{1})=\quad\sum^{U^{P}+U^{P^{\prime}},U^{M^{0}}+U^{P_{1}}\dots}_{\mathclap{\begin{subarray}{c}\{U^{M^{0}}=|U^{P}-U^{P^{\prime}}|,\,U^{M}=|U^{M^{0}}-U^{P_{1}}|,\,U^{F}\dots\}\\ U^{P}_{3}+U^{P^{\prime}}_{3}=U_{3}^{M_{0}},\,U^{M_{0}}_{3}+U_{3}^{P_{1}}=U_{3}^{M}\\ U_{3}^{M}+U_{3}^{B_{1}}=U_{3}^{F},\,U_{3}^{F}-U_{3}^{B_{b}}=U_{3}^{\mathcal{H}}\end{subarray}}}C^{\{U^{P}_{3}U^{P^{\prime}}_{3}U_{3}^{M_{0}}\}}_{U^{P}U^{P^{\prime}}U^{M_{0}}}C^{U_{3}^{M_{0}}\,U_{3}^{P_{1}}\,U_{3}^{M}}_{U^{M^{0}}\,U^{P_{1}}\,U^{M}}$ (6)
$\times C^{U_{3}^{M}\,U_{3}^{B_{1}}\,U_{3}^{F}}_{U^{M}\,U^{B_{1}}\,U^{F}}C^{U_{3}^{F}\,(-U_{3}^{B_{b}})\,U_{3}^{\mathcal{H}}}_{U^{F}\,U^{B_{b}}\,U^{\mathcal{H}}}A^{U^{M},U^{F},\pm}(U^{\mathcal{H}})$
where $A^{i,j,\pm}(U^{\mathcal{H}})$ denotes the $U$-spin subamplitudes and $C^{U_{3}^{a}U_{3}^{b}U_{3}^{c}}_{U^{a}U^{b}U^{c}}$ represents the $SU(2)$ Clebsch–Gordan coefficients. The $\pm$ label in $A^{i,j,\pm}(U^{\mathcal{H}})$ distinguishes the $U$-spin doublet to which the final state light baryon belongs, and in general $A^{i,j,+}(U^{\mathcal{H}})$ and $A^{i,j,-}(U^{\mathcal{H}})$ are unrelated. The subamplitudes $A^{i,j,\pm}(U^{\mathcal{H}})$ in the most general case also involve unknown dynamical coefficients, for example, in the form of Wilson coefficients and CKM factors, that only get fixed when a particular form of the effective Hamiltonian is assumed. Therefore, apriori $A^{i,j,\pm}(U^{\mathcal{H}})$ are all independent of each other. We now proceed to a discussion of the lowest order effective Hamiltonian [41, 42, 43] mediating charmless bottom baryon $\Delta|S|=1$ and $\Delta S=0$ transitions. Both parts of the effective Hamiltonian is composed of the operators $O_{1}$,…, $O_{10}$ and can be written as:
$\mathcal{H}_{\text{eff}}=\frac{4G_{F}}{\sqrt{2}}\Big[\lambda^{(s)}_{u}\Big(C_{1}(O^{(u)}_{1}-O^{(c)}_{1})+C_{2}(O^{(u)}_{2}-O^{(c)}_{2})\Big)-\lambda^{(s)}_{t}\sum_{i=1,2}C_{i}O^{(c)}_{i}-\lambda^{(s)}_{t}\sum_{i=3}^{10}C_{i}O_{i}^{(s)}$
$+\lambda^{(d)}_{u}\Big(C_{1}(O^{(u)}_{1}-O^{(c)}_{1})+C_{2}(O^{(u)}_{2}-O^{(c)}_{2})\Big)-\lambda^{(d)}_{t}\sum_{i=1,2}C_{i}O^{(c)}_{i}-\lambda^{(d)}_{t}\sum_{i=3}^{10}C_{i}O_{i}^{(d)}\Big],$ (7)
where $V_{ub}V_{us}^{*}=\lambda_{u}^{s}$, $V_{ub}V_{ud}^{*}=\lambda_{u}^{d}$, $V_{tb}V_{ts}^{*}=\lambda_{t}^{s}$, $V_{tb}V_{td}^{*}=\lambda_{t}^{d}$ are the CKM elements and $C_{i}$ s are the Wilson coefficients. $O_{1}$ and $O_{2}$ are the “tree” operators:
$O_{1}^{(u)}$ $=(\overline{u}^{i}_{L}\gamma^{\mu}b^{j}_{L})(\overline{s}^{j}_{L}\gamma_{\mu}u^{i}_{L})$
$O_{1}^{(c)}$ $=(\overline{c}^{i}_{L}\gamma^{\mu}b^{j}_{L})(\overline{s}^{j}_{L}\gamma_{\mu}c^{i}_{L})$
$O_{2}^{(u)}$ $=(\overline{u}^{i}_{L}\gamma^{\mu}b^{i}_{L})(\overline{s}^{j}_{L}\gamma_{\mu}u^{j}_{L})$
$O_{2}^{(c)}$ $=(\overline{c}^{i}_{L}\gamma^{\mu}b^{i}_{L})(\overline{s}^{j}_{L}\gamma_{\mu}c^{j}_{L}).$ (8)
$O_{3}$, …, $O_{6}$ are the “gluonic penguin” operators:
$O_{3}^{(s)}$ $=(\overline{s}^{i}_{L}\gamma^{\mu}b^{i}_{L})\sum_{q=u,d,s}(\overline{q}^{j}_{L}\gamma_{\mu}q^{j}_{L})$
$O_{4}^{(s)}$ $=(\overline{s}^{i}_{L}\gamma^{\mu}b^{j}_{L})\sum_{q=u,d,s}(\overline{q}^{j}_{L}\gamma_{\mu}q^{i}_{L})$
$O_{5}^{(s)}$ $=(\overline{s}^{i}_{L}\gamma^{\mu}b^{i}_{L})\sum_{q=u,d,s}(\overline{q}^{j}_{R}\gamma_{\mu}q^{j}_{R})$
$O_{6}^{(s)}$ $=(\overline{s}^{i}_{L}\gamma^{\mu}b^{j}_{L})\sum_{q=u,d,s}(\overline{q}^{j}_{R}\gamma_{\mu}q^{i}_{R}).$ (9)
and finally $O_{7}$, …, $O_{10}$ are the four “Electroweak penguin” operators:
$O_{7}^{(s)}$ $=\frac{3}{2}(\overline{s}^{i}_{L}\gamma^{\mu}b^{i}_{L})\sum_{q=u,d,s}e_{q}(\overline{q}^{j}_{R}\gamma_{\mu}q^{j}_{R}),$
$O_{8}^{(s)}$ $=\frac{3}{2}(\overline{s}^{i}_{L}\gamma^{\mu}b^{j}_{L})\sum_{q=u,d,s}e_{q}(\overline{q}^{j}_{R}\gamma_{\mu}q^{i}_{R}),$ (10)
$O_{9}^{(s)}$ $=\frac{3}{2}(\overline{s}^{i}_{L}\gamma^{\mu}b^{i}_{L})\sum_{q=u,d,s}e_{q}(\overline{q}^{j}_{L}\gamma_{\mu}q^{j}_{L}),$
$O_{10}^{(s)}$ $=\frac{3}{2}(\overline{s}^{i}_{L}\gamma^{\mu}b^{j}_{L})\sum_{q=u,d,s}e_{q}(\overline{q}^{j}_{L}\gamma_{\mu}q^{i}_{L}).$ (11)
Restricting ourselves to a discussion of $U$-spin of these operators, we find that the tree operators $O_{1}$ and $O_{2}$ transform as a $U$-spin doublet with the up component corresponding to $b\to d$ transition ($\Delta S=0$) while the down component corresponds to $b\to s$ transition ($\Delta|S|=1$) transition. The relevant CKM elements which form part of the dynamical coefficients are given, respectively, as $V_{ub}V_{ud}^{*}$ and $V_{ub}V_{us}^{*}$. Since the combination $d\overline{d}+s\overline{s}$ transforms as a $U$-spin singlet, the complete set of QCD and Electroweak penguin four-Fermi operators ($O_{3}\dots O_{10}$) transform as distinct $U$-spin doublets. Interestingly, unlike isospin, which is explicitly broken by electromagnetic interactions, leading to isospin-violating effects in the electroweak penguin operators, no such complication arises in the case of $U$-spin, since the $d$ and $s$ quarks carry the same electric charge. Therefore, to the leading order in $U$-spin, the dim-6 SM effective Hamiltonian only induces a doublet in Eq (6). We observe that certain higher dimensional $U$-spin representations present in the general Hamiltonian have vanishing dynamical coefficients (See Appendix (A)) corresponding to non-existent $U$-spin subamplitudes.
III Amplitude Relations
Possible four body all-electrically charged final states originating from charmless bottom baryon decays are listed in Table (1).
$\Delta|S|=1$ modes $\Delta S=0$ modes $\Delta|S|=1$ modes $\Delta S=0$ modes
$\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+}$ $\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+}$ $\Xi^{0}_{b}\to p^{+}K^{-}K^{-}\pi^{+}$ $\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+}$
$\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-}$ $\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-}$ $\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}K^{-}$ $\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}\pi^{-}$
$\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+}$ $\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+}$ $\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}\pi^{+}$ $\Lambda^{0}_{b}\to p^{+}\pi^{-}K^{-}K^{+}$
$\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-}$ $\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-}$ $\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-}$ $\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}\pi^{-}$
$\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+}$ $\Xi^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+}$ $\Xi^{0}_{b}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-}$ $\Lambda^{0}_{b}\to\Sigma^{-}K^{+}K^{+}K^{-}$
$\Lambda^{0}_{b}\to\Xi^{-}K^{+}\pi^{+}\pi^{-}$ $\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}K^{+}K^{-}$ $\Xi^{0}_{b}\to\Sigma^{+}K^{-}K^{-}K^{+}$ $\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+}$
$\Lambda^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+}$ $\Xi^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+}$ $\Lambda^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-}$ $\Xi^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-}$
Table 1: Charmless four body decays of bottom baryons involving all charged final states
We choose to represent the $\Delta S=0$ and $\Delta|S|=1$ decay amplitudes and the complete set of $U$-spin subamplitudes in terms of column matrices $\mathcal{A}$ and $A$ respectively such that there is coefficient matrix ($\mathbb{C}$) that links both basis through the equation,
$\mathcal{A}=\mathbb{C}A$ (12)
The explicit form the matrix $\mathbb{C}$ is given in Appendix (A). We notice that there are in general 14 independent $U$-spin subamplitudes describing all 28 decays. Alternatively, these 28 decay modes can be grouped based on the charge of the final state baryon. In that case, 14 decay modes with a positively charged baryon are described in terms of 7 $A^{i,j,+}(U^{\mathcal{H}})$ subamplitudes while the remaining 14 decay modes with a negatively charged baryon are described in terms of 7 $A^{i,j,-}(U^{\mathcal{H}})$ subamplitudes. At this stage, no simple amplitude relation between only two decay modes can be established. Now assuming the SM Hamiltonian expressed in Eq (II), which contributes to the $U$-spin subamplitudes $A^{\frac{1}{2},0,\pm}(U^{\mathcal{H}}=\frac{1}{2}),\,A^{\frac{1}{2},1,\pm}(U^{\mathcal{H}}=\frac{1}{2})$ and $A^{\frac{3}{2},1,\pm}(U^{\mathcal{H}}=\frac{1}{2})$ but not to $A^{\frac{1}{2},1,\pm}(U^{\mathcal{H}}=\frac{3}{2}),\,A^{\frac{3}{2},1,\pm}(U^{\mathcal{H}}=\frac{3}{2}),\,A^{\frac{3}{2},2,\pm}(U^{\mathcal{H}}=\frac{3}{2}),\,A^{\frac{3}{2},2,\pm}(U^{\mathcal{H}}=\frac{5}{2})$ for both the tree and penguin operators, the number of independent $U$-spin subamplitudes can be reduced to 6, three for each class of decays with a positively or negatively charged baryon in the final state. We can further decompose the decay amplitudes in terms of tree ($\mathcal{A}_{T}$) and penguin ($\mathcal{A}_{P}$) parts by factoring out the CKM elements $\lambda^{s,d}_{u,t}$ from the decay amplitude and write,
$\mathcal{A}^{(s,d)}=\lambda_{u}^{s,d}\mathcal{A}_{T}+\lambda^{s,d}_{t}\mathcal{A}_{P}$ (13)
where $(s,d)$ denote the $\Delta|S|=1$, $\Delta S=0$ process respectively.
$B_{b}\to B_{1}^{+}PP^{{}^{\prime}}P_{1}$ modes $B_{b}\to B_{1}^{-}PP^{{}^{\prime}}P_{1}$ modes $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}K^{-}K^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}K^{+}K^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}\pi^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}\pi^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to p^{+}\pi^{-}K^{-}K^{+})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})$ $\mathcal{A}_{T(P)}(\Lambda^{0}_{b}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})=\mathcal{A}_{T(P)}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}K^{-})$
Table 2: Amplitude $U$-spin relations between $\Delta|S|=1$ and $\Delta S=0$ four body bottom baryon decays, assuming the unbroken SM Hamiltonian. The subscript $T$ and $P$ denotes the tree and penguin part of the full decay amplitude respectively.
We obtain the following amplitude relations connecting the $\Delta|S|=1$ and $\Delta S=0$ sectors given in Table (2). $A_{CP}$ is defined subsequently as,
$A_{CP}=$ $\frac{\Gamma(B_{b}\to B_{1}PP^{\prime}P_{1})-\Gamma(\overline{B_{b}}\to\overline{B_{1}}\overline{P}\overline{P^{\prime}}\overline{P_{1}})}{\Gamma(B_{b}\to B_{1}PP^{\prime}P_{1})+\Gamma(\overline{B_{b}}\to\overline{B_{1}}\overline{P}\overline{P^{\prime}}\overline{P_{1}})}$
$=$  $\frac{-\,4\textbf{J}\times\text{Im}\Big[\mathcal{A}^{*}_{T}(B_{b}\to B_{1}PP^{\prime}P_{1})\mathcal{A}_{P}(B_{b}\to B_{1}PP^{\prime}P_{1})\Big]}{|\mathcal{A}(B_{b}\to B_{1}PP^{\prime}P_{1})|^{2}+|\mathcal{A}(\overline{B_{b}}\to\overline{B_{1}}\overline{P}\overline{P^{\prime}}\overline{P_{1}})|^{2}},$ (14)
where J is the Jarlskog invariant given by $\text{Im}(V_{ub}V_{ud}^{*}V_{tb}^{*}V_{td})=-\text{Im}(V_{ub}V_{us}^{*}V_{tb}^{*}V_{ts})=\textbf{J}$. Neglecting any $U$-spin breaking effects from phase space factors, we can relate the $CP$-asymmetry in $\Delta|S|=1$ and $\Delta S=0$ decay modes using Eq (15),
$\frac{A_{CP}(B_{bi}\to B_{1}^{a}P^{b}P^{{}^{\prime}c}P_{1}^{d})|_{\Delta S}}{A_{CP}(B_{bj}\to B_{1}^{e}P^{f}P^{{}^{\prime}g}P_{1}^{h})|_{\Delta S^{\prime}}}\!\simeq-\frac{\tau_{\mathcal{B}_{bi}}}{\tau_{\mathcal{B}_{bj}}}\frac{Br(B_{bj}\to B_{1}^{e}P^{f}P^{{}^{\prime}g}P_{1}^{h})}{Br(B_{bi}\to B_{1}^{a}P^{b}P^{{}^{\prime}c}P_{1}^{d}))},\qquad\Delta S\neq\Delta S^{\prime}$ (15)
where $i$, $j$, and a,ea,\,e b,fb,\,f, c,gc,\,g, d,hd,\,h indices correspond to the various baryon and mesons in the initial and final states and the negative sign accounts for the relative sign between the Jarlskog factor in $\Delta|S|=1$ and $\Delta S=0$ decay modes. These $A_{CP}$-relations can be read off from Table (2) by using Eq (15). As reported by LHCb, the $\Lambda^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+}$ decay as well as other four-body bottom baryon decays occur primarily through hadronic resonances that individually decay into two or three final-state particles. The contribution from multiple resonances results in substantial variation in associated strong phases and relative strengths of tree ($\mathcal{A}_{T}$) and penguin amplitudes ($\mathcal{A}_{P}$) leading to differences in the $CP$-asymmetry across the final-state phase space of the bottom baryon decay. Therefore, it is reasonable to investigate local $CP$-asymmetries near resonant regions in such decays, as will be explored in the following. If all the final-state mesons originate from a system with $U$-spin = 1/2, as observed in one of the four phase-space regions of the decay $\Lambda^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+}$ [1], the $A^{\frac{3}{2},1,\pm}(\frac{1}{2})$ $U$-spin subamplitudes can be dropped. This simplification allows us to describe decays in Table (2) in terms of 4 $U$-spin subamplitudes, two for each class of decays with a positively or negatively charged baryon in the final state. This leads to additional $CP$-asymmetry relations that are given by,
$A_{CP}(\Lambda^{0}_{b}\to p^{+}(K^{-}K^{-}K^{+}))$ $\simeq A_{CP}(\Lambda^{0}_{b}\to p^{+}(\pi^{-}K^{-}\pi^{+})),$
$=-A_{CP}(\Xi_{b}^{0}\to\Sigma^{+}(\pi^{-}\pi^{-}\pi^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})}{Br(\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})}$
$=-\frac{1}{2}A_{CP}(\Xi_{b}^{0}\to\Sigma^{+}(\pi^{-}K^{-}K^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}K^{+})}{Br(\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})}$ (16a)
$A_{CP}(\Lambda^{0}_{b}\to p^{+}(\pi^{-}\pi^{-}\pi^{+}))$ $\simeq A_{CP}(\Lambda^{0}_{b}\to p^{+}(\pi^{-}K^{-}K^{+}))$
$=-A_{CP}(\Xi_{b}^{0}\to\Sigma^{+}(K^{-}K^{-}K^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})}{Br(\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+})}$
$=-\frac{1}{2}A_{CP}(\Xi_{b}^{0}\to\Sigma^{+}(\pi^{-}K^{-}\pi^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})}{Br(\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+})}$ (16b)
$A_{CP}(\Lambda^{0}_{b}\to\Sigma^{+}(\pi^{-}\pi^{-}\pi^{+}))$ $\simeq A_{CP}(\Lambda^{0}_{b}\to\Sigma^{+}(\pi^{-}K^{-}K^{+}))$
$=-A_{CP}(\Xi_{b}^{0}\to p^{+}(K^{-}K^{-}K^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})}{Br(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})}$
$=-\frac{1}{2}A_{CP}(\Xi_{b}^{0}\to p^{+}(\pi^{-}K^{-}\pi^{+}))\frac{\tau_{\Lambda^{0}_{b}}}{\tau_{\Xi^{0}_{b}}}\frac{Br(\Xi_{b}^{0}\to p^{+}\pi^{-}K^{-}\pi^{+})}{Br(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})}$ (16c)
Measurements of the following branching ratios of $\Lambda^{0}_{b}$-baryon and $\Xi^{0}_{b}$-baryon are available from LHCb [44],
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})=(5.1\pm 0.5)\times 10^{-5}$
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}K^{+})=(4.1\pm 0.6)\times 10^{-6}$
$Br(\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})=(1.27\pm 0.13)\times 10^{-5}$ (17)
$Br(\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+})=(2.12\pm 0.21)\times 10^{-5}$
$Br(\Xi^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})\times\frac{F(b\to\Xi_{b}^{0})}{F(b\to\Lambda^{0}_{b})}=(1.9\pm 0.4)\times 10^{-6}$
$Br(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}\pi^{+})\times\frac{F(b\to\Xi_{b}^{0})}{F(b\to\Lambda^{0}_{b})}=(1.73\pm 0.31)\times 10^{-6}$
$Br(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\times\frac{F(b\to\Xi_{b}^{0})}{F(b\to\Lambda^{0}_{b})}=(1.8\pm 1.0)\times 10^{-7}$
where $F(b\to B_{b})$ denotes the fragmentation fraction of $b$-quarks into a $B_{b}$-baryon. From the available local $CP$-asymmetry measurements in bottom baryon decays [13, 1], we identify that $CP$-asymmetry in $\Lambda_{b}^{0}\to p^{+}K_{1}^{-}(1410)$ quasi-two body decay [13],
$A_{CP}(\Lambda_{b}^{0}\to p^{+}(\pi^{-}K^{-}\pi^{+}))=(4.7\pm 3.5\pm 0.8)\%,$ (18)
can be translated into local $CP$-asymmetry in $\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}\pi^{+}$ transition since $K_{1}^{-}(1410)$ decays strongly to $\pi^{+}K^{-}\pi^{-}$ [44]. Moreover, $K_{1}^{-}(1410)\to K^{+}K^{-}K^{-}$ strong decay is also possible [45] and therefore based on Eq (16a) we expect that local $CP$-asymmetry in $\Lambda^{0}_{b}\to p^{+}K^{-}K^{-}K^{+}$ transition is,
$A_{CP}(\Lambda_{b}^{0}\to p^{+}(K^{-}K^{-}K^{+})_{K_{1}})=A_{CP}(\Lambda_{b}^{0}\to p^{+}(\pi^{-}K^{-}\pi^{+})_{K_{1}})\simeq(4.7\pm 3.6)\%$ (19)
We observe that the predicted value of $A_{CP}(\Lambda_{b}^{0}\to p^{+}(\pi^{-}K^{-}\pi^{+})_{K_{1}})$ has the same sign as the recently measured local $A_{CP}$ [1]
$A_{CP}(\Lambda_{b}^{0}\to p^{+}(\pi^{-}K^{-}\pi^{+}))|_{m_{\pi^{+}K^{-}\pi^{-}}<\,2.0\,\text{GeV}}=(2.0\pm 1.2\pm 0.3)\%,$ (20)
although the central values differ by roughly a factor of two. Similarly, we focus on $CP$-violation in the quasi-two body $\Lambda^{0}_{b}\to p^{+}a_{1}^{-}(1260)$ decay where we note that $a_{1}^{-}(1260)$ subsequently can decay strongly to either $\pi^{+}\pi^{-}\pi^{-}$ or $K^{+}\pi^{-}K^{-}$ [44]. We use the measured $CP$-asymmetry in $\Lambda^{0}_{b}\to p^{+}a_{1}^{-}(1260)$-decay [13],
$A_{CP}(\Lambda^{0}_{b}\to p^{+}a_{1}^{-}(1260))=(-1.5\pm 4.2\pm 0.6)\%$ (21)
and with the help of Eq (16b) also predict local $CP$-asymmetry in $\Lambda^{0}_{b}\to p^{+}\pi^{-}K^{-}K^{+}$ and $\Lambda^{0}_{b}\to p^{+}\pi^{-}\pi^{-}\pi^{+}$,
$A_{CP}(\Lambda^{0}_{b}\to p^{+}(\pi^{-}\pi^{-}\pi^{+})_{a_{1}})=A_{CP}(\Lambda^{0}_{b}\to p^{+}(\pi^{-}K^{-}K^{+})_{a_{1}})\simeq(-1.5\pm 4.2)\%.$ (22)
Once the branching fractions and local $CP$ asymmetries of additional $\Lambda^{0}_{b}$ and $\Xi^{0}_{b}$ decay modes, such as $\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+}$ and $\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}K^{-}K^{+}$ become available, they can be used to express the local-$CP$ asymmetries in related $\Xi^{0}_{b}$ decay modes: $\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+}$ and $\Xi^{0}_{b}\to p^{+}\pi^{-}K^{-}\pi^{+}$ with the help of Eq (16c). The $CP$-asymmetry relations inferred from Table (2) can also be tested by experiments in near future. We now discuss the effect of first order breaking of $U$-spin originating form unequal masses of the $d$ an $s$ quark. $U$-spin breaking is introduced by a triplet under $U$-spin ($U_{\text{brk}}=1,U^{3}_{\text{brk}}=0$) and by taking the tensor product with the unbroken Hamiltonian that transforms as a doublet. The effect of this $U$-spin breaking is a correction to the unbroken Hamiltonian given by,
$\mathcal{H}_{\text{brk}}^{\Delta|S|=0,1}\propto\mathcal{\epsilon}(1,0)\otimes(\frac{1}{2},\pm\frac{1}{2})=\epsilon\Big(\sqrt{\frac{2}{3}}(\frac{3}{2},\pm\frac{1}{2})\mp\sqrt{\frac{1}{3}}(\frac{1}{2},\pm\frac{1}{2})\Big)$ (23)
This induces non-zero dynamical coefficients for the $U$-spin subamplitudes $A^{\frac{1}{2},1,\pm}(\frac{3}{2}),\,A^{\frac{3}{2},1,\pm}(\frac{3}{2}),$ $A^{\frac{3}{2},2,\pm}(\frac{3}{2})$ while $A^{\frac{1}{2},0}(\frac{1}{2}),\,A^{\frac{1}{2},1}(\frac{1}{2}),\,A^{\frac{3}{2},1}(\frac{1}{2})$ pick up $U$-spin breaking corrections. Collectively, we can parametrize these $U$-spin breaking amplitudes as,
$A_{\text{brk}}^{\frac{1}{2},0,\pm}(\frac{1}{2})=\frac{\epsilon_{1}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},0,\pm}(\frac{1}{2}),\qquad A_{\text{brk}}^{\frac{1}{2},1,\pm}(\frac{1}{2})=\frac{\epsilon_{2}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},1,\pm}(\frac{1}{2})$
$A_{\text{brk}}^{\frac{1}{2},1,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{3}^{\pm}}{\sqrt{3}}A^{\frac{1}{2},1,\pm}(\frac{3}{2}),\qquad A_{\text{brk}}^{\frac{3}{2},1,\pm}(\frac{1}{2})=\frac{\epsilon_{4}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},1,\pm}(\frac{1}{2}),$ (24)
$A_{\text{brk}}^{\frac{3}{2},1,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{5}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},1,\pm}(\frac{3}{2}),\qquad A_{\text{brk}}^{\frac{3}{2},2,\pm}(\frac{3}{2})=\frac{\sqrt{2}\epsilon_{6}^{\pm}}{\sqrt{3}}A^{\frac{3}{2},2,\pm}(\frac{3}{2})$
The combined result of leading order unbroken Hamiltonian and first order $U$-spin breaking contribution to the decay amplitudes is given in Appendix (B). The first order $U$-spin breaking corrections have the same structures but opposite signs for the decay modes related by $U$-spin as found in [29]. This increases the rank of the coefficient matrices to 12 and the leading order relations between $U$-spin conjugate amplitude pairs no longer hold. Naturally, the leading order $CP$-asymmetry relations evident from Table (2) and Eq (16) also break down once $U$-spin breaking effects are considered. Any reduction among the set of $U$-spin subamplitudes would require knowledge of their relative magnitudes, which lies beyond the scope of the present work.
IV Conclusion
We have investigated charmless bottom baryon decays into fully charged four-body final states within the realm of $U$-spin symmetry. Bearing in mind that significant $U$-spin breaking is observed in $B$-meson decays, we set up a framework to analyze bottom baryon decays where $U$-spin symmetry can be arbitrarily broken. We obtained amplitude relations between decay modes and based on those relations predicted $CP$ asymmetry relations between a number of decay modes. By making the additional assumption that all final-state mesons emerge from a $U$-spin doublet, we demonstrated that this extended framework allows for numerical predictions of localized $CP$ asymmetries in specific regions of phase space, which hold to a good accuracy in practice. The framework also predicts the relative signs of the $CP$ asymmetries, an important feature that can be obscured in global $CP$ asymmetry measurements due to averaging over the entire phase space of the decay. While $U$-spin breaking effect spoils the relations between observables, the predictions offer a roadmap for probing $CP$ violation in bottom baryon decays and provide a systematic approach to test the validity of $U$-spin symmetry in future experimental measurements.
Appendix A
$U$-spin amplitude decomposition of $B_{b}\to B_{1}^{+}PP^{\prime}P_{1}$ decay modes for general Hamiltonian:
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}K^{+})\\ \end{pmatrix}=\left(\begin{array}[]{ccccccc}\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&-\frac{1}{\sqrt{5}}\\ -\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&-\frac{1}{\sqrt{10}}\\ 0&\frac{2}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&0&0&-\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&-\frac{1}{\sqrt{10}}\\ 0&\frac{\sqrt{2}}{3}&-\frac{1}{3}&\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&-\frac{1}{\sqrt{5}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&-\frac{2}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&-\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&-\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&-\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{3}{2})\\ A^{\frac{3}{2},1,+}(\frac{1}{2})\\ A^{\frac{3}{2},1,+}(\frac{3}{2})\\ A^{\frac{3}{2},2,+}(\frac{3}{2})\\ A^{\frac{3}{2},2,+}(\frac{5}{2})\end{pmatrix}$ (39)
$U$-spin amplitude decomposition of $B_{b}\to B_{1}^{-}PP^{\prime}P_{1}$ decay modes for general Hamiltonian:
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}K^{+}\pi^{-})\\ \end{pmatrix}=\left(\begin{array}[]{ccccccc}-\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&-\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&-\frac{1}{\sqrt{5}}\\ 0&\frac{2}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{\sqrt{2}}{3}&-\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ 0&\frac{\sqrt{2}}{3}&-\frac{1}{3}&\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&-\frac{1}{\sqrt{5}}\\ 0&0&0&-\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&-\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{\sqrt{15}}&\frac{1}{\sqrt{10}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&-\sqrt{\frac{2}{15}}&\frac{1}{\sqrt{5}}\\ 0&-\frac{2}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&\frac{1}{6}&-\frac{\sqrt{\frac{3}{5}}}{2}&-\frac{1}{\sqrt{10}}\\ \frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&-\frac{1}{3}&\frac{\sqrt{2}}{3}&\sqrt{\frac{2}{15}}&-\frac{1}{\sqrt{5}}\\ -\frac{1}{\sqrt{3}}&\frac{1}{3}&-\frac{\sqrt{2}}{3}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&\frac{1}{\sqrt{15}}&-\frac{1}{\sqrt{10}}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&\sqrt{\frac{3}{10}}&\frac{1}{\sqrt{5}}\\ 0&0&0&\frac{1}{\sqrt{2}}&\frac{1}{2}&\frac{\sqrt{\frac{3}{5}}}{2}&\frac{1}{\sqrt{10}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{3}{2})\\ A^{\frac{3}{2},1,-}(\frac{1}{2})\\ A^{\frac{3}{2},1,-}(\frac{3}{2})\\ A^{\frac{3}{2},2,-}(\frac{3}{2})\\ A^{\frac{3}{2},2,-}(\frac{5}{2})\end{pmatrix}$ (54)
$U$-spin amplitude decomposition of $B_{b}\to B_{1}PP^{\prime}P_{1}$ decay modes using the SM Hamiltonian,
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}K^{+}\pi^{-})\\ \end{pmatrix}=\mathcal{C}_{T(P)}\left(\begin{array}[]{cccccc}\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}\\ -\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}\\ 0&0&0&0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}\\ 0&\frac{2}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ \frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}\\ -\frac{1}{\sqrt{6}}&-\frac{1}{3\sqrt{2}}&\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&-\frac{1}{3}&-\frac{1}{3\sqrt{2}}\\ 0&0&-\frac{1}{\sqrt{2}}&0&0&0\\ 0&0&0&0&\frac{\sqrt{2}}{3}&\frac{1}{3}\\ 0&\frac{\sqrt{2}}{3}&\frac{1}{3}&0&0&0\\ 0&0&0&0&0&-\frac{1}{\sqrt{2}}\\ -\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}\\ \frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}\\ 0&0&0&0&-\frac{2}{3}&\frac{1}{3\sqrt{2}}\\ 0&-\frac{2}{3}&\frac{1}{3\sqrt{2}}&0&0&0\\ -\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}&0&0&0\\ 0&0&0&\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}\\ \frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}&-\frac{1}{3}&0&0&0\\ 0&0&0&-\frac{1}{\sqrt{3}}&\frac{1}{3}&\frac{1}{3\sqrt{2}}\\ 0&0&\frac{1}{\sqrt{2}}&0&0&0\\ 0&0&0&0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}\\ 0&-\frac{\sqrt{2}}{3}&-\frac{1}{3}&0&0&0\\ 0&0&0&0&0&\frac{1}{\sqrt{2}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{1}{2})\\ A^{\frac{3}{2},1,+}(\frac{1}{2})\\ A^{\frac{1}{2},0,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{1}{2})\\ A^{\frac{3}{2},1,-}(\frac{1}{2})\\ \end{pmatrix}$ (83)
where $\mathcal{C}_{T}$ and $\mathcal{C}_{P}$ are the Wilson Coefficients of tree operators ($O_{1},\,O_{2}$) and penguin operators ($O_{3}\dots O_{10}$) respectively.
Appendix B
$U$-spin amplitude decomposition of $B_{b}\to B_{1}PP^{\prime}P_{1}$ decay modes incorporating first order $U$-spin breaking to the SM Hamiltonian,
$\begin{pmatrix}\mathcal{A}(\Lambda^{0}_{b}\to p^{+}K^{-}\pi^{-}\pi^{+}\\ \mathcal{A}(\Lambda^{0}_{b}\to\Xi^{-}K^{+}K^{+}K^{-}\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Lambda^{0}_{b}\to\Sigma^{-}K^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}K^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}K^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}\pi^{+}K^{-})\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{+}K^{-}\pi^{-}K^{+}\\ \mathcal{A}(\Xi^{0}_{b}\to\Sigma^{-}\pi^{+}\pi^{+}\pi^{-}\\ \mathcal{A}(\Xi^{0}_{b}\to p^{+}K^{-}K^{-}K^{+})\\ \mathcal{A}(\Xi^{0}_{b}\to\Xi^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}K^{-})\\ \mathcal{A}(\Xi_{b}^{0}\to p^{+}K^{-}\pi^{-}\pi^{+})\\ \mathcal{A}(\Xi_{b}^{0}\to\Xi^{-}K^{+}K^{+}K^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{+}\pi^{-}\pi^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Sigma^{-}\pi^{+}K^{+}\pi^{-})\\ \mathcal{A}(\Lambda_{b}^{0}\to p^{+}\pi^{-}K^{-}K^{+})\\ \mathcal{A}(\Lambda_{b}^{0}\to\Xi^{-}K^{+}K^{+}\pi^{-})\\ \end{pmatrix}=\left(\begin{array}[]{cccccccccccc}\frac{\left(\epsilon_{1}\right){}^{+}}{3\sqrt{2}}+\frac{1}{\sqrt{6}}&-\frac{\left(\epsilon_{2}\right){}^{+}}{3\sqrt{6}}-\frac{1}{3\sqrt{2}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{+}&\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{3}}+\frac{1}{3}&\frac{2\left(\epsilon_{5}\right){}^{+}}{3\sqrt{3}}&-\frac{2\left(\epsilon_{6}\right){}^{+}}{3\sqrt{5}}&0&0&0&0&0&0\\ 0&0&0&0&0&0&-\frac{1}{3}\left(\epsilon_{1}\right){}^{-}-\frac{1}{\sqrt{3}}&-\frac{\left(\epsilon_{2}\right){}^{-}}{3\sqrt{3}}-\frac{1}{3}&-\frac{2\left(\epsilon_{3}\right){}^{-}}{3\sqrt{3}}&-\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{6}}-\frac{1}{3\sqrt{2}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{5}\right){}^{-}&-\frac{1}{3}\sqrt{\frac{2}{5}}\left(\epsilon_{6}\right){}^{-}\\ -\frac{1}{3}\left(\epsilon_{1}\right){}^{+}-\frac{1}{\sqrt{3}}&-\frac{\left(\epsilon_{2}\right){}^{+}}{3\sqrt{3}}-\frac{1}{3}&-\frac{2\left(\epsilon_{3}\right){}^{+}}{3\sqrt{3}}&-\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{6}}-\frac{1}{3\sqrt{2}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{5}\right){}^{+}&-\frac{1}{3}\sqrt{\frac{2}{5}}\left(\epsilon_{6}\right){}^{+}&0&0&0&0&0&0\\ 0&0&0&0&0&0&\frac{\left(\epsilon_{1}\right){}^{-}}{3\sqrt{2}}+\frac{1}{\sqrt{6}}&-\frac{\left(\epsilon_{2}\right){}^{-}}{3\sqrt{6}}-\frac{1}{3\sqrt{2}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{-}&\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{3}}+\frac{1}{3}&\frac{2\left(\epsilon_{5}\right){}^{-}}{3\sqrt{3}}&-\frac{2\left(\epsilon_{6}\right){}^{-}}{3\sqrt{5}}\\ 0&0&0&0&0&0&0&\frac{2\left(\epsilon_{2}\right){}^{-}}{3\sqrt{3}}+\frac{2}{3}&-\frac{2\left(\epsilon_{3}\right){}^{-}}{3\sqrt{3}}&-\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{6}}-\frac{1}{3\sqrt{2}}&\frac{\left(\epsilon_{5}\right){}^{-}}{3\sqrt{6}}&-\frac{\left(\epsilon_{6}\right){}^{-}}{\sqrt{10}}\\ 0&\frac{2\left(\epsilon_{2}\right){}^{+}}{3\sqrt{3}}+\frac{2}{3}&-\frac{2\left(\epsilon_{3}\right){}^{+}}{3\sqrt{3}}&-\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{6}}-\frac{1}{3\sqrt{2}}&\frac{\left(\epsilon_{5}\right){}^{+}}{3\sqrt{6}}&-\frac{\left(\epsilon_{6}\right){}^{+}}{\sqrt{10}}&0&0&0&0&0&0\\ \frac{\left(\epsilon_{1}\right){}^{+}}{3}+\frac{1}{\sqrt{3}}&-\frac{\left(\epsilon_{2}\right){}^{+}}{3\sqrt{3}}-\frac{1}{3}&-\frac{2\left(\epsilon_{3}\right){}^{+}}{3\sqrt{3}}&-\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{6}}-\frac{1}{3\sqrt{2}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{5}\right){}^{+}&\frac{1}{3}\sqrt{\frac{2}{5}}\left(\epsilon_{6}\right){}^{+}&0&0&0&0&0&0\\ 0&0&0&0&0&0&-\frac{\left(\epsilon_{1}\right){}^{-}}{3\sqrt{2}}-\frac{1}{\sqrt{6}}&-\frac{\left(\epsilon_{2}\right){}^{-}}{3\sqrt{6}}-\frac{1}{3\sqrt{2}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{-}&\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{3}}+\frac{1}{3}&\frac{2\left(\epsilon_{5}\right){}^{-}}{3\sqrt{3}}&\frac{2\left(\epsilon_{6}\right){}^{-}}{3\sqrt{5}}\\ -\frac{\left(\epsilon_{1}\right){}^{+}}{3\sqrt{2}}-\frac{1}{\sqrt{6}}&-\frac{\left(\epsilon_{2}\right){}^{+}}{3\sqrt{6}}-\frac{1}{3\sqrt{2}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{+}&\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{3}}+\frac{1}{3}&\frac{2\left(\epsilon_{5}\right){}^{+}}{3\sqrt{3}}&\frac{2\left(\epsilon_{6}\right){}^{+}}{3\sqrt{5}}&0&0&0&0&0&0\\ 0&0&0&0&0&0&\frac{\left(\epsilon_{1}\right){}^{-}}{3}+\frac{1}{\sqrt{3}}&-\frac{\left(\epsilon_{2}\right){}^{-}}{3\sqrt{3}}-\frac{1}{3}&-\frac{2\left(\epsilon_{3}\right){}^{-}}{3\sqrt{3}}&-\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{6}}-\frac{1}{3\sqrt{2}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{5}\right){}^{-}&\frac{1}{3}\sqrt{\frac{2}{5}}\left(\epsilon_{6}\right){}^{-}\\ 0&0&0&-\frac{\left(\epsilon_{4}\right){}^{+}}{\sqrt{6}}-\frac{1}{\sqrt{2}}&\frac{\left(\epsilon_{5}\right){}^{+}}{\sqrt{6}}&\frac{\left(\epsilon_{6}\right){}^{+}}{\sqrt{10}}&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{2}\right){}^{-}+\frac{\sqrt{2}}{3}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{-}&\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{3}}+\frac{1}{3}&-\frac{\left(\epsilon_{5}\right){}^{-}}{3\sqrt{3}}&\frac{\left(\epsilon_{6}\right){}^{-}}{\sqrt{5}}\\ 0&\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{2}\right){}^{+}+\frac{\sqrt{2}}{3}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{+}&\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{3}}+\frac{1}{3}&-\frac{\left(\epsilon_{5}\right){}^{+}}{3\sqrt{3}}&\frac{\left(\epsilon_{6}\right){}^{+}}{\sqrt{5}}&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&-\frac{\left(\epsilon_{4}\right){}^{-}}{\sqrt{6}}-\frac{1}{\sqrt{2}}&\frac{\left(\epsilon_{5}\right){}^{-}}{\sqrt{6}}&\frac{\left(\epsilon_{6}\right){}^{-}}{\sqrt{10}}\\ \frac{\left(\epsilon_{1}\right){}^{+}}{3\sqrt{2}}-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}-\frac{\left(\epsilon_{2}\right){}^{+}}{3\sqrt{6}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{+}&\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{3}}-\frac{1}{3}&\frac{2\left(\epsilon_{5}\right){}^{+}}{3\sqrt{3}}&-\frac{2\left(\epsilon_{6}\right){}^{+}}{3\sqrt{5}}&0&0&0&0&0&0\\ 0&0&0&0&0&0&\frac{1}{\sqrt{3}}-\frac{\left(\epsilon_{1}\right){}^{-}}{3}&\frac{1}{3}-\frac{\left(\epsilon_{2}\right){}^{-}}{3\sqrt{3}}&-\frac{2\left(\epsilon_{3}\right){}^{-}}{3\sqrt{3}}&\frac{1}{3\sqrt{2}}-\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{6}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{5}\right){}^{-}&-\frac{1}{\sqrt{15}}\\ \frac{1}{\sqrt{3}}-\frac{\left(\epsilon_{1}\right){}^{+}}{3}&\frac{1}{3}-\frac{\left(\epsilon_{2}\right){}^{+}}{3\sqrt{3}}&-\frac{2\left(\epsilon_{3}\right){}^{+}}{3\sqrt{3}}&\frac{1}{3\sqrt{2}}-\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{6}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{5}\right){}^{+}&-\frac{1}{3}\sqrt{\frac{2}{5}}\left(\epsilon_{6}\right){}^{+}&0&0&0&0&0&0\\ 0&0&0&0&0&0&\frac{\left(\epsilon_{1}\right){}^{-}}{3\sqrt{2}}-\frac{1}{\sqrt{6}}&\frac{1}{3\sqrt{2}}-\frac{\left(\epsilon_{2}\right){}^{-}}{3\sqrt{6}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{-}&\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{3}}-\frac{1}{3}&\frac{2\left(\epsilon_{5}\right){}^{-}}{3\sqrt{3}}&-\frac{2\left(\epsilon_{6}\right){}^{-}}{3\sqrt{5}}\\ 0&0&0&0&0&0&0&\frac{2\left(\epsilon_{2}\right){}^{-}}{3\sqrt{3}}-\frac{2}{3}&-\frac{2\left(\epsilon_{3}\right){}^{-}}{3\sqrt{3}}&\frac{1}{3\sqrt{2}}-\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{6}}&\frac{\left(\epsilon_{5}\right){}^{-}}{3\sqrt{6}}&-\frac{\left(\epsilon_{6}\right){}^{-}}{\sqrt{10}}\\ 0&\frac{2\left(\epsilon_{2}\right){}^{+}}{3\sqrt{3}}-\frac{2}{3}&-\frac{2\left(\epsilon_{3}\right){}^{+}}{3\sqrt{3}}&\frac{1}{3\sqrt{2}}-\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{6}}&\frac{\left(\epsilon_{5}\right){}^{+}}{3\sqrt{6}}&-\frac{\left(\epsilon_{6}\right){}^{+}}{\sqrt{10}}&0&0&0&0&0&0\\ \frac{\left(\epsilon_{1}\right){}^{+}}{3}-\frac{1}{\sqrt{3}}&\frac{1}{3}-\frac{\left(\epsilon_{2}\right){}^{+}}{3\sqrt{3}}&-\frac{2\left(\epsilon_{3}\right){}^{+}}{3\sqrt{3}}&\frac{1}{3\sqrt{2}}-\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{6}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{5}\right){}^{+}&\frac{1}{3}\sqrt{\frac{2}{5}}\left(\epsilon_{6}\right){}^{+}&0&0&0&0&0&0\\ 0&0&0&0&0&0&\frac{1}{\sqrt{6}}-\frac{\left(\epsilon_{1}\right){}^{-}}{3\sqrt{2}}&\frac{1}{3\sqrt{2}}-\frac{\left(\epsilon_{2}\right){}^{-}}{3\sqrt{6}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{-}&\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{3}}-\frac{1}{3}&\frac{2\left(\epsilon_{5}\right){}^{-}}{3\sqrt{3}}&\frac{2\left(\epsilon_{6}\right){}^{-}}{3\sqrt{5}}\\ \frac{1}{\sqrt{6}}-\frac{\left(\epsilon_{1}\right){}^{+}}{3\sqrt{2}}&\frac{1}{3\sqrt{2}}-\frac{\left(\epsilon_{2}\right){}^{+}}{3\sqrt{6}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{+}&\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{3}}-\frac{1}{3}&\frac{2\left(\epsilon_{5}\right){}^{+}}{3\sqrt{3}}&\frac{2\left(\epsilon_{6}\right){}^{+}}{3\sqrt{5}}&0&0&0&0&0&0\\ 0&0&0&0&0&0&\frac{\left(\epsilon_{1}\right){}^{-}}{3}-\frac{1}{\sqrt{3}}&\frac{1}{3}-\frac{\left(\epsilon_{2}\right){}^{-}}{3\sqrt{3}}&-\frac{2\left(\epsilon_{3}\right){}^{-}}{3\sqrt{3}}&\frac{1}{3\sqrt{2}}-\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{6}}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{5}\right){}^{-}&\frac{1}{3}\sqrt{\frac{2}{5}}\left(\epsilon_{6}\right){}^{-}\\ 0&0&0&\frac{1}{\sqrt{2}}-\frac{\left(\epsilon_{4}\right){}^{+}}{\sqrt{6}}&\frac{\left(\epsilon_{5}\right){}^{+}}{\sqrt{6}}&\frac{\left(\epsilon_{6}\right){}^{+}}{\sqrt{10}}&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{2}\right){}^{-}-\frac{\sqrt{2}}{3}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{-}&\frac{\left(\epsilon_{4}\right){}^{-}}{3\sqrt{3}}-\frac{1}{3}&-\frac{\left(\epsilon_{5}\right){}^{-}}{3\sqrt{3}}&\frac{\left(\epsilon_{6}\right){}^{-}}{\sqrt{5}}\\ 0&\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{2}\right){}^{+}-\frac{\sqrt{2}}{3}&-\frac{1}{3}\sqrt{\frac{2}{3}}\left(\epsilon_{3}\right){}^{+}&\frac{\left(\epsilon_{4}\right){}^{+}}{3\sqrt{3}}-\frac{1}{3}&-\frac{\left(\epsilon_{5}\right){}^{+}}{3\sqrt{3}}&\frac{\left(\epsilon_{6}\right){}^{+}}{\sqrt{5}}&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&\frac{1}{\sqrt{2}}-\frac{\left(\epsilon_{4}\right){}^{-}}{\sqrt{6}}&\frac{\left(\epsilon_{5}\right){}^{-}}{\sqrt{6}}&\frac{\left(\epsilon_{6}\right){}^{-}}{\sqrt{10}}\\ \end{array}\right)\begin{pmatrix}A^{\frac{1}{2},0,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,+}(\frac{3}{2})\\ A^{\frac{3}{2},1,+}(\frac{1}{2})\\ A^{\frac{3}{2},1,+}(\frac{3}{2})\\ A^{\frac{3}{2},2,+}(\frac{3}{2})\\ A^{\frac{1}{2},0,+}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{1}{2})\\ A^{\frac{1}{2},1,-}(\frac{3}{2})\\ A^{\frac{3}{2},1,-}(\frac{1}{2})\\ A^{\frac{3}{2},1,-}(\frac{3}{2})\\ A^{\frac{3}{2},2,-}(\frac{3}{2})\\ \end{pmatrix}$